What is the probability of flipping a coin 7 times and receiving all heads?
The probability of each coin coming up heads is 1/2 .
The probability of all 7 coming up heads is (1/2 x 1/2 x 1/2 x 1/2 x 1/2 x 1/2 x 1/2) = 1/128
That's 0.78 percent (rounded)
(Dr. Atticus was right.)
The expected number is 12/7 = 1.29 approx.
What is the probability of getting exactly 5 boys and 5 girls if you have 10 kids?
Here the sample space is(s)=10, =>mod(s)=10
a=be the event of getting exactly 5 boys =>mod(a)=5
b=be the event of getting exactly 5 girls =>mod(b)=5
thus,
p(a)=mod(a)/mod(s)=5/10=1/2
p(b)=mod(b)/mod(s)=5/10=1/2
p(5 boys and 5 girls)=p(a)*p(b)=1/2*1/2=1/4
Bala has 5, Ramu has 7
It is (4/8)*(6/8) = 3/8
How many different ways can you arrange 15 cars in 12 parking spaces?
Any 12 from 15 in any order = 15 x 14 x 13 x 12 x 11 x 10 x 9 x 8 x 7 x 6 x 5 x 4 = 217,945,728,000 different ways!
This seems ridiculously high but...
The first space can be any one of fifteen, the second, any one of fourteen and so on until the twelfth can be any one of the remaining four.
8
hundredths
Construction of two random variables such that their correlation coefficient equals -1?
y = cos ( x ).
z = cos ( x + 180 degrees ).
Two 6 sided dice are rolled What is the probability the sum of the numbers is 12 or 4?
For a sum = 12, there's only one (6 + 6), 1/36
For a sum = 4, there's three: 2/36 (1 + 3) + 1/36 (2 + 2) = 3/36
What is the probability of getting 5 or 6 on a single throw?
There are 6 outcomes on a die roll (1,2,3,4,5,6); 5 & 6 constitutes 2 of them so the probability is 2/6 or 1/3.
I am not sure but I think 3/36
The difference between the theoretical and empirical probability?
The term empirical means "based on observation or experiment." An empirical probability is generally, but not always, given with a number indicating the possible percent error (e.g. 80+/-3%). A theoretical probability, however, is one that is calculatedbased on theory, i.e., without running any experiments.
Since there is no theory that will calculate the probability that an area will experience an earthquake within a given time frame, the 90% figure is an empirical probability, presumably based on data of major earthquakes in the San Francisco area over past years.
20 marbles total, 10 of which are blue
so, odds of drawing blue are 10/20=.5=50%
it depends on how many time you throw the dice. the increase number of your throw, the higher the probability to get 8 that one die is showing a 5. that what my opinion.